OpenStudy (anonymous):

Help Please!!! 1.A kid run around a field which shapes is rectangle with the length 16 Meter. if he run around the field two times,so he already managed the distance of 108 Meter. a. create a mathematic model from that,and finish it. b. what is broad field such? i really appreciate anyone who help me right now. thank you :)

OpenStudy (anonymous):

the field is 16 meters long and 2 laps around it is 108 meters. That means that the perimeter is 54 meters. 2 sides are 16 meters which means that the other 2 sides have to be 11 meters. As for the model of the field, it would be a rectangle with the dimensions of 16x11. I could be wrong but that is my understanding of your question

OpenStudy (anonymous):

i see.. but what i asked is in a. and b. .. but i aprreciate your answer,though. thank you very much :D

OpenStudy (anonymous):

your welcome hope i helped you in some way

OpenStudy (anonymous):

yeah,you helped me find how to search the broad of the field :D again,thank you very much :)

OpenStudy (anonymous):

good luck with the rest of it

OpenStudy (anonymous):

haha,yeah. i'll try then :)

OpenStudy (aravindg):

mathematic model ? you mean frame an equation ?

OpenStudy (aravindg):

@jiehanthelearner ?

OpenStudy (anonymous):

yeah,some sort. i mean,a mathematical problem from that topic

OpenStudy (aravindg):

ok its a rectangle so perimeter will be 2(l+b) where l is length and b is breadth . Now the kid made 2 runs that will be 2 X 2(l+b)=4(l+b) It is said the kid managed 108 meter when doing this .Hence we can equate this to 2 times perimeter 4(l+b)=108 Also given length=16 m substituting 4(16+b)=108 ^^This I suppose is the mathematical model of the question Now to find breadth just open the brackets and solve for b I hope that helps :)

OpenStudy (anonymous):

thanks!! :D

OpenStudy (aravindg):

yw :)